# A Global Minimum Tax for Large Firms Only:

### v5

@@ 5 Conclusion
This paper has analyzed the effects of a global minimum tax (GMT) that is confined to large multinational enterprises (MNEs), thus leaving at least 10% of the global multinational tax base outside its scope. Using a simple model with profit shifting by heterogeneous MNEs, we have shown that this partial coverage of the GMT gives rise to a sequence of tax competition equilibria between a non-haven and a set of symmetric tax haven countries. In particular, introducing a low GMT rate that still binds the tax havens induces them to commit to a single GMT rate for all MNEs and thus results in higher welfare and tax revenues in both sets of countries. However, a further increase in the GMT rate leads first the tax havens, and then the non-haven, to split their tax rate for large and small MNEs, creating a conflict of interest between the two groups of countries. A similar conflict also arises from broadening the coverage of MNEs subject to the GMT.
The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, the most likely regime is one where the non-haven sets a uniform tax rate above the GMT for all MNEs, whereas the haven countries split their tax rate and undercut the GMT for small MNEs (Regime 2). This finding corresponds to the observation that tax havens such as Ireland, Liechtenstein, Bermuda, or Bulgaria maintain their regular corporate tax rates below the level of 15%, but top up the taxation of large multinationals to the level of the GMT. In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the tax havens lose, resulting in a positive net gain globally.
Our main objective in this paper has been to understand the implications of an incomplete coverage of the GMT. We have done so in a highly stylized model, in order to develop some sharp intuitions. With the insights gained, it would be fruitful to extend our basic model to incorporate further important aspects of the global minimum tax and the responses of multinational firms. A first extension would allow MNEs to adjust their real pre-tax profits in response to changes in tax rates, for example, by modifying their investment decision. A second extension would incorporate a splitting response of multinationals seeking to benefit from potentially lower tax rates on small MNEs. A third extension would consider the substance-based income exclusion, i.e., additional tax deductions for MNEs that increase investment and employment. We leave these extensions for further research.
Appendix A: Proof of Propositions
A.1. Proof of Proposition 1
Regime 0 can be an equilibrium as long as the haven’s unconstrained rate is greater than the GMT rate, $t_{M}<t_{h}^{0}=t_{M}^{0}\equiv\delta(\lambda-1)/[H(3\lambda-1)]$. A new equilibrium regime emerges when $t_{M}\geq t_{M}^{0}$. Since haven countries are all symmetric, we are concerned with the representative haven country $h$ in what follows.
@@ A.1. Proof of Proposition 1
$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\qquad G_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
When the non-haven sets a single non-GMT rate and the haven splits its tax rates, the equilibrium tax rates are
@@
$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\qquad G_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
@@
where the strict equality holds at $t_{M}\in\{t_{M}^{0},t_{M}^{1}\}$. Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate, and the haven sets a single GMT rate, given by (A.1). Table A1 shows the relevant payoffs in Regimes 1 and 2, where the GMT rate is not binding for the non-haven.
Non-haven Haven
single GMT rate split tax rate
single non-GMT rate $G_{n}^{1},\quad G_{h}^{1}$ $G_{n}^{2},\quad G_{h}^{2}$
single GMT rate
split tax rate
Table A1: Payoff matrix for Regimes 1 ($t_{M}\in[t_{M}^{1},t_{M}^{2}]$) and 2 ($t_{M}\in(t_{M}^{1},t_{M}^{2}]$)
Notes: In Regimes 1 and 2, the GMT is binding only for the haven. The potential regimes where the non-haven does choose a single GMT rate or split tax rates are irrelevant because, whenever possible, it always chooses a single non-GMT rate that solves the unconstrained maximization problem.
@@
Since (A.3) becomes negative for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, the haven splits its tax rates in response to a non-haven’s single non-GMT rate. The non-haven is still not bound by the GMT and continues to chooses its optimal unconstrained tax rate. Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate and the haven splits its tax rates, with tax rates given by (A.2).
@@
When the non-haven sets a single GMT rate and the haven splits its tax rates, the equilibrium tax rates and payoffs are
$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\qquad G_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
@@
$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\qquad G_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
@@
which is always fulfilled in Regime 3. The strict equality holds at $t_{M}=t_{M}^{3}$ and it holds that $t_{M}>t_{M}^{2}>\frac{2\delta(\lambda-1)}{H(3\lambda-1)}$ for any $\lambda>1/3$. Therefore, the tax-competition equilibrium is that the non-haven sets a single GMT rate, and the haven splits its tax rates, as given in (A.4).
@@
Since (A.6) becomes negative for $t_{M}\in(t_{M}^{3},1]$, the non-haven splits its tax rates in response to the havens’ split rates. Therefore, the tax-competition equilibrium is that both the non-haven and the haven split their tax rates, as given in (A.5).
Non-haven Haven
single GMT rate split tax rate
single non-GMT rate
single GMT rate $G_{n}^{3},\quad G_{h}^{3}$
split tax rate $G_{n}^{4},\quad G_{h}^{4}$
Table A2: Payoff matrix for Regimes 3 ($t_{M}\in(t_{M}^{2},t_{M}^{3}]$) and 4 ($t_{M}\in(t_{M}^{3},1]$)
Notes: In Regimes 3 and 4, the GMT is binding for both the non-haven and the haven. Therefore, the regimes where the non-haven chooses a single-non GMT rate are irrelevant. The regimes where the haven sets a single GMT are also irrelevant because, by doing so, the haven would obtain non-positive welfare/tax revenues.
A.2. Decentralized decision of tax havens
@@
In our main model, we have assumed that tax havens collectively decide on whether to set a single GMT rate, or to split their tax rate. We here examine an alternative, decentralized decision process in the first stage of our game where each single tax haven decides whether to split or not, given the other havens’ decisions. To simplify this analysis, we assume that countries can split their tax rate only if they are bound by the GMT. This reflects the fact that preferential tax regimes for specific tax bases are not allowed under the OECD’s action plan to fight base erosion.
@@
Suppose then that $t_{M}\in[t_{M}^{0},t_{M}^{2}]$. Out of the total of $H$ tax havens, there are $H_{M}$ havens that commit to a single GMT rate while the remaining $H_{S}=H-H_{M}$ havens split their tax rate. When the non-haven sets a single non-GMT rate and a single haven $h$ sets a uniform GMT rate, the equilibrium tax rates are
$t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\qquad t_{h}^{1}=t_{M},$ (A.7)
which is the counterpart of (A.1). Let $G_{h}^{1}(H_{M})$ be the haven’s equilibrium payoff with commitment.
If the single haven $h$ splits its tax rate instead, whereas the non-haven sets a uniform non-GMT rate, the equilibrium tax rates are
@@
$t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\qquad t_{h}^{2}=\begin{cases}t_{n}(H_{M})/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},$ (A.8)
which is the counterpart of (A.2). Let $G_{h}^{2}(H_{M})$ be the haven’s equilibrium payoff under splitting.
We then ask if a single haven has an incentive to choose a uniform GMT rate given that all the other $H-1$ havens split their tax rates. A single haven $h$ has a unilateral incentive to do so, if the following holds:
$\Delta G_{h}^{12}(H_{M}=1)$ $\equiv G_{h}^{1}(H_{M}=1)-G_{h}^{2}(H_{M}=0)$
$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta}{\delta H^{2}[\lambda(3+\phi)-(1+\phi)]^{2}[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]}\geq 0,$ (A.9)
where $\Theta\equiv A-Bt_{M},\quad A\equiv\delta(\lambda-1)[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]>0,$
$B\equiv H\phi(\lambda-1)[\lambda(3H-5)-(H-3)]+H(3\lambda-1)[\lambda(3H-1)-(H-1)]>0,$
$\text{and}\qquad Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0\ \ \text{or}\ \ t_{M}\geq\frac{\delta(\lambda-1)}{H(3\lambda-1)}\equiv t_{M}^{0}.$
It is straightforward to show that $\Theta>0$ holds at $t_{M}^{0}$ and hence, by continuity, also for $t_{M}$ marginally above $t_{M}^{0}$. Therefore, there must be a positive range of $t_{M}$, with $t_{M}^{0}<t_{M}\leq t_{M}^{a}\equiv A/B$ for which the inequality (A.9) holds for a haven $h$. By symmetry, this argument applies to all havens.
For a Regime 1 equilibrium with commitment to occur, a further condition is that no single haven $h$ must have an incentive to split, given that all the other $H_{M}-1$ havens commit to a single GMT rate. A single haven $h$ has no unilateral incentive to split its tax rate, if the following holds:
$\Delta G_{h}^{12}(H_{M}=H)$ $\equiv G_{h}^{1}(H_{M}=H)-G_{h}^{2}(H_{M}=H-1)$
$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta^{\prime}}{\delta H(2\lambda-1)[\phi(\lambda-1)+\lambda(4H-1)-(2H-1)]^{2}}\geq 0,$ (A.10)
where $\Theta^{\prime}\equiv C-Dt_{M},\quad C\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
$D\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
and $t_{M}\geq t_{M}^{0}$ holds again. The equivalent condition is $t_{M}\leq t_{M}^{b}\equiv C/D$. We can immediately see $\Delta G_{h}^{12}(H_{M}=H)<0$ for $t_{M}\in(t_{M}^{b},t_{M}^{2}]$ and check $t_{M}^{a}<t_{M}^{b}<t_{M}^{1}<t_{M}^{2}$:
$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{\Theta^{\prime\prime}[(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}]}>0,$
$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}}{\Theta^{\prime\prime}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0,$
$\Theta^{\prime\prime}\equiv 3(3H-1)\lambda^{2}-2(3H-2)\lambda+H-1+\phi(\lambda-1)[\lambda(3H-5)-(H-3)]>0.$
@@
We can conclude that for $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, all the havens set a single GMT rate and the non-haven sets a single non-GMT rate; for $t_{M}\in(t_{M}^{a},t_{M}^{2}]$, all the havens split their tax rate and the non-haven sets a single non-GMT rate. The former regime corresponds to Regime 1 and the latter one to Regime 2 in the text. The only difference is that the range of Regime 1 is smaller when havens make their first stage decision in a decentralized way, implying that the havens’ commitment is less likely. This is intuitive because the commitment of a single haven does not induce the non-haven to increase its tax rate by as much as the simultaneous commitment of many havens. Finally, in the remaining range of the GMT, $t_{M}\in(t_{M}^{2},1]$, the difference between collective vs. decentralized decisions of havens plays no role any more. In either setting, the tax havens always have an incentive to split their tax rate, whether the non-haven sets a single GMT rate (Regime 3), or splits its tax rate as well (Regime 4).
A.3. Proofs of Propositions 3 and 4

### v6

@@ 4 Conclusion
This paper has analysed the effects of a global minimum tax (GMT) confined to large multinational enterprises (MNEs), leaving about 10% of the global tax base outside its scope. This partial-coverage GMT generates a sequence of tax competition equilibria between a non-haven and tax havens. In particular, introducing a low GMT rate that binds the havens induces them to commit to the single GMT rate. However, a further increase in the GMT rate leads first the havens, and then the non-haven, to split their tax rates for large and small MNEs.
The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, the most likely regime is one where the non-haven sets a uniform tax rate above the GMT for all MNEs, whereas the havens split their tax rates and undercut the GMT for small MNEs. This finding corresponds to the observation that tax havens such as Ireland and Switzerland maintain their base corporate tax rates below 15%, but top up taxes on large MNEs to the level of the GMT. In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the havens lose, resulting in a positive net gain globally.
Appendix
Appendix 1: Proof of Proposition 1
Regime 0 can be an equilibrium as long as the haven’s unconstrained rate is greater than the GMT rate, $t_{M}<t_{h}^{0}=t_{M}^{0}\equiv\delta(\lambda-1)/[H(3\lambda-1)]$. A new equilibrium regime emerges when $t_{M}\geq t_{M}^{0}$. Since haven countries are all symmetric, we are concerned with a representative haven country $h$ in what follows.
@@
$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\hskip 18.49988ptG_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
When the non-haven sets a single non-GMT rate and the haven splits its tax rate, the equilibrium tax rates are
@@
$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\hskip 18.49988ptG_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
@@
where the strict equality holds at $t_{M}\in\{t_{M}^{0},t_{M}^{1}\}$. Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate, and the haven sets a single GMT rate, given by (A.1).
@@
Since (A.3) becomes negative for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, the haven splits its tax rate in response to a non-haven’s single non-GMT rate. The non-haven is still not bound by the GMT and continues to choose its unconstrained optimal tax rate. Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate and the haven splits its tax rate (see (A.2)).
@@
When the non-haven sets a single GMT rate and the haven splits its tax rate, the equilibrium tax rates and payoffs are
$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\hskip 18.49988ptt_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\hskip 18.49988ptG_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
@@
$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\hskip 18.49988ptt_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\hskip 18.49988ptG_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
@@
which is always fulfilled in Regime 3. The strict equality holds at $t_{M}=t_{M}^{3}$ and it holds that $t_{M}>t_{M}^{2}>\frac{2\delta(\lambda-1)}{H(3\lambda-1)}$ for any $\lambda>1/3$. Therefore, the tax-competition equilibrium is that the non-haven sets a single GMT rate, and the haven splits its tax rate, as given in (A.4).
@@
Since (A.6) becomes negative for $t_{M}\in(t_{M}^{3},1]$, the non-haven splits its tax rate in response to the haven’s split tax rates. Therefore, the tax-competition equilibrium is that both the non-haven and the haven split their tax rates. ∎
Supplemental Appendix of
‘A Global Minimum Tax for Large Firms Only:
Implications for Tax Competition’
Andreas Haufler [^1] Seminar for Economic Policy, LMU Munich, Akademiestr. 1, 80799 Munich, Germany. Email: Andreas.Haufler@econ.lmu.de.     Hayato Kato [^2] Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan. Email: hayato.kato@econ.osaka-u.ac.jp
October, 2025
Contents
Appendix A MNEs Subject to the GMT
We select from the Orbis database MNEs satisfying the following conditions:
  • Non-missing revenues (‘Operating revenue (Turnover)’) and pre-tax profits (‘P/L before tax’).
Column (a) of Table A.1 reports the pre-tax profits and the total number of MNEs in each year. Column (b) restricts the sample to MNEs subject to the GMT, i.e., those with annual revenues of at least 750 million EUR in two or more of the past four years. Using columns (a) and (b), column (c) shows the share of MNEs subject to the GMT in terms of the pre-tax profits and the number of MNEs.
Table A.1. MNEs in the Orbis Database.
(a) All MNEs (b) MNEs $\geq$ 750mEUR (c) Share == B/A
2018 2320 8656 2066 2333 0.89 0.27
2019 2430 8416 2216 2437 0.91 0.29
2020 1985 7920 1793 2228 0.90 0.28
2021 3564 8803 3239 2627 0.91 0.30
Source: Orbis database, own calculations.
Note: Pre-tax profits are in billion EUR.
Appendix B Decentralized Decision-Making by Tax Havens
@@
In the text, we assumed that a group of tax-haven countries collectively decide whether to adopt a single GMT rate or to split their tax rate. Here, we consider an alternative, decentralized decision-making process in the first stage of our game, where each tax haven individually decides whether to split, given the decisions of the others. In the following, we focus on $t_{M}\in[t_{M}^{0},t_{M}^{2}]$ and impose two assumptions. First, countries can split their tax rates only if they are subject to the GMT. This assumption reflects the fact that preferential tax regimes for specific tax bases are not permitted under the OECD’s action plan to address base erosion and profit shifting (OECD, 2019). Second, each haven forms the same belief about how the other havens and the non-haven choose their tax rates. This allows us to focus on the problem of a single haven.
@@
Out of the total $H$ tax havens, $H_{M}$ havens commit to a single GMT rate, while the remaining $H-H_{M}$ havens split their tax rates. When the non-haven sets a single non-GMT rate and a haven $h$ sets a single GMT rate, their equilibrium tax rates are respectively
$\begin{split}&t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\\ &t_{h}^{1}(H_{M})=t_{M},\end{split}$
Let $G_{h}^{1}(H_{M})$ denote the equilibrium payoff of the haven $h$ with commitment. We note that as in the text, setting a single non-GMT rate is the non-haven’s dominant strategy because if $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, it is never constrained by the GMT: $t_{n}(H_{M})\leq t_{n}(H)=t_{n}^{2}\leq t_{M}$.
If instead the haven $h$ splits its tax rate while the other havens and the non-haven keep their rates unchanged, the number of havens that maintain commitment falls to $H_{M}-1$. The equilibrium tax rates of the non-haven and the haven $h$ are respectively
@@
$\begin{split}&t_{n}(H_{M}-1)=\frac{2(\lambda-1)\left[t_{M}\left\{H\phi+(H_{M}-1)(1-\phi)\right\}+\delta\right]}{(\lambda-1)[\phi(H-H_{M}+1)+H+H_{M}-1]+2H\lambda},\\ &t_{h}^{2}(H_{M}-1)=\begin{cases}t_{n}(H_{M}-1)/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\end{split}$ (B.1)
Let $G_{h}^{2}(H_{M}-1)$ be the equilibrium payoff of the haven $h$ under splitting.
The difference between the haven $h$’s payoffs in the two cases is
$\Delta G_{h}^{12}(H_{M})\equiv G_{h}^{1}(H_{M})-G_{h}^{2}(H_{M}-1)$
$\ \ =\frac{(1-\phi)\Pi[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta(H_{M})}{\delta\left[\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})\right]\left[(\lambda-1)\left\{\phi(H-H_{M}+1)+H+H_{M}-1\right\}+2H\lambda\right]^{2}},$
where
$\Theta(H_{M})\equiv AH_{M}+B,$
$A\equiv(\lambda-1)(1-\phi)\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right],$
$B\equiv H[\phi(\lambda-1)+3\lambda-1]\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right]-2t_{M}(\lambda-1)^{2}(1-\phi).$
If $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$, the haven $h$ always sets a single GMT rate, regardless of how many other havens it believes set a single GMT rate. Under our assumption of symmetric belief formation, all havens end up setting a single GMT rate, and the equilibrium coincides with Regime 1 in the text. Conversely, if $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$, all havens end up splitting their tax rates, and the equilibrium coincides with Regime 2 in the text.
Since the denominator of $\Delta G_{h}^{12}(H_{M})$ is positive and $Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0$ for $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, the sign of $\Delta G_{h}^{12}(H_{M})$ is determined by that of $\Theta(H_{M})$. Considering that $\Theta(H_{M})$ is linear in $H_{M}$, the conditions for $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$ reduce to the following two conditions:
$\Theta(1)=A\cdot 1+B=C-Dt_{M}\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{a}\equiv C/D,$
$\Theta(H)=AH+B=2(E-Ft_{M})\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{b}\equiv E/F,$
where
$C\equiv\delta(\lambda-1)[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]>0,$
$D\equiv H\phi(\lambda-1)[\lambda(3H-5)-(H-3)]+H(3\lambda-1)[\lambda(3H-1)-(H-1)]>0,$
$E\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
$F\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
We can check $t_{M}^{0}<t_{M}^{a}<t_{M}^{b}<t_{M}^{1}$:
$t_{M}^{a}-t_{M}^{0}=\frac{2\delta(\lambda-1)^{2}[\phi(\lambda-1)+3\lambda-1]}{D(3\lambda-1)}>0,$
$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{DF}>0,$
$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}[\phi(H+1)(\lambda-1)+\lambda(3H-1)-(H-1)]}{F[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0.$
We can also confirm $\Theta(H_{M})=AH_{M}+B$ increases with $H_{M}$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$, or equivalently $A>0$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$. Since $A>0$ if $t_{M}<t_{M}^{c}\equiv\delta(\lambda-1)/[H(3\lambda-1)-2(\lambda-1)]$, it suffices to show $t_{M}^{b}<t_{M}^{c}$:
$t_{M}^{c}-t_{M}^{b}=\frac{\delta(\lambda-1)[\lambda^{2}+(H-1)(2\lambda-1)-\phi(\lambda-1)^{2}]}{F[H(3\lambda-2)-(H-1)]}>0,$
which is indeed positive.
Using these results, we can thus conclude that (i) if $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$; (ii) if $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$; and (iii) if $t_{M}\in(t_{M}^{a},t_{M}^{b}]$, there exists a positive integer $H_{M}^{*}\in\{2,\dots,H\}$ such that $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,\dots,H_{M}^{*}-1\}$ and $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{H_{M}^{*},\dots,H\}$.
In cases (i) and (ii), the equilibrium is unique. In case (iii), there can be at most three equilibria. First, if each haven believes that fewer than $H_{M}^{*}$ havens set a single GMT rate, all havens end up splitting their tax rates, i.e., $H_{M}=0$. Second, if each haven believes that more than $H_{M}^{*}-1$ havens set a single GMT rate, all havens end up doing so, i.e., $H_{M}=H$. Third, if $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, then $H_{M}=H_{M}^{*}$ havens set a single GMT rate while the remaining $H-H_{M}^{*}$ havens split their tax rates, consistent with their beliefs.
The discussion is summarized as follows. If $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, the non-haven chooses a single non-GMT rate, and each haven chooses a single GMT rate, as in Regime 1 in the text (Proposition 1(i)). If $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, the non-haven sets a single non-GMT rate, and each haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs, as in Regime 2 in the text (Proposition 1(ii)). If $t_{M}\in[t_{M}^{a},t_{M}^{b})$, both Regimes 1 and 2 can be equilibria and either may emerge. If $t_{M}\in[t_{M}^{a},t_{M}^{b})$ and in addition $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, the non-haven sets a single non-GMT rate ($t_{n}(H_{M}^{*})$ in (B.1)), and there are $H_{M}^{*}$ havens setting a single GMT rate $t_{M}$ and $H-H_{M}^{*}$ havens splitting their tax rates ($t_{h}^{2}(H_{M}^{*})$ in (B.1)).
@@
A clear difference from Proposition 1 is that the range where only Regime 1 occurs is smaller ($t_{M}^{0},t_{M}^{a})\subset[t_{M}^{0},t_{M}^{1})$) when havens make their first stage decision in a decentralized manner, implying that they are less likely to commit to a single GMT rate. This is intuitive because the commitment of a haven induces the non-haven to raise its tax rate less than the simultaneous commitment of many havens.
Appendix C Welfare Effects of the Introduction of GMT
The following proposition characterizes the welfare effects of introducing a binding GMT rate, compared with the unconstrained equilibrium (Regime 0).
Proposition C.1
(GMT introduction)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text, and assume the GMT coverage rate is sufficiently high, such that $\phi\in(3/4,1)$. Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
  • (i)
    a rise in welfare in the non-haven for all Regimes 1 to 4.
  • (ii)
    a rise in welfare in the (representative) haven for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
  • (iii)
    a rise in world welfare for all Regimes 1 to 4.
These results hold for any value of the marginal value of public funds (MVPF), $\lambda\in(1,\infty)$. In the limit as $\lambda\to\infty$, countries care only about tax revenues. Hence, qualitatively identical results apply to tax revenues.
Proof: In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.